OpenStudy (anonymous):

a right triangle has one leg that is 8inches. the hypotenuse is 17 inches. find the length of the other leg

OpenStudy (anonymous):

its 15 inches

OpenStudy (anonymous):

Thanks a bunch=)

OpenStudy (anonymous):

8^2+b^2=17^2 squareroot(17^2-8^2)=b

OpenStudy (anonymous):

b=15